EX 2ok when soka Full td JJ X E B 1 -A 1 су t 3 By 번 am 3m Aynusing force method
Answer #1

Method ormotion of consistent Using deformati to find reactions Since, structure has no fle=0 hz load & is in one plane Ar reof super position & Using Method standard for mulce to calculate final deflection which will finally be zero. at B4L Due to u3 Jom/2 . Sc₂ до x 12 3 СІ х2 2EZ 433 4o 1 О. we to By , 9 By xg3 243 By 31 S : В) 9- By x 9* 3 E7 2 І 4 R6 By су “ to Cy д пfinal deflection Сp - о 8 : 8,1 8в, * 6», ;dczo - (A 1202 8. f 29 30 Agent чөс : EZ 1 EZ 2 43 В, + +16 с. : чі веѕ Se: €, ғS,=> Ay = 26-294 KW EMA =0 -MA + 50 20x 9 x 2x9 -117, 912 49 3 of Box12 25.794x is to =) IM ² 51.882 kNm A